OpenStudy (gabylovesyou):

Choose the slope-intercept equation of the line that passes through the point (-5, -1) and is perpendicular to y = 5/2x + 2.

OpenStudy (gabylovesyou):

y = 2/5x + 1 y = -5/2x - 27/2 y = -2/5x - 3 y = 5/2x + 23/2

OpenStudy (gabylovesyou):

@cwrw238

OpenStudy (gabylovesyou):

@ash2326

OpenStudy (ash2326):

@Gabylovesyou did you try this?

OpenStudy (gabylovesyou):

yes but i dont get how to ...

OpenStudy (ash2326):

Ok. Standard slope and intercept form equation of a line \[y=mx+c\] where m= slope c= y-intercept here we have a line \[y=\frac 52 x+2\] What is its slope?

OpenStudy (gabylovesyou):

to find the slope i have to.... ?

OpenStudy (ash2326):

compare standard equation and the given equation.

OpenStudy (ash2326):

do you get this?

OpenStudy (gabylovesyou):

mmm kinda

OpenStudy (ash2326):

Read my second post again and compare to find the slope of the line

OpenStudy (gabylovesyou):

mmmmmmmmm can u do it step by step???

OpenStudy (ash2326):

Should I given you an example?

OpenStudy (gabylovesyou):

yes please

OpenStudy (ash2326):

suppose I have a line \[y=5x+6\] I need to find the slope I'll compare it with the standard equation \[y=mx+c\] I see m=5 and c=6 so slope =5 and y- intercept=6 do you get this?

OpenStudy (gabylovesyou):

ok... mmm so slope is 5/2 ?

OpenStudy (ash2326):

Yes

OpenStudy (ash2326):

if two lines are perpendicular then \[m1\times m2=-1\] m1 and m2 are their slope respecticely here m1= 5/2, what's m2?

OpenStudy (gabylovesyou):

2?

OpenStudy (ash2326):

\[\frac 52 \times m2 =-1\] \[m2=???\]

OpenStudy (gabylovesyou):

,,,, - 5/4 ?

OpenStudy (ash2326):

@Gabylovesyou How you'd solve this? \[\frac 1 2\times x=1\]

OpenStudy (gabylovesyou):

x = 2

OpenStudy (ash2326):

solve this also the same way \[\frac 52 \times m_2=-1\]

OpenStudy (gabylovesyou):

idk how too